# Autonomous Navigation and Station-Keeping on Near-Rectilinear Halo Orbits

### v1

@@ 6 Controller Experiments
Prior to running the full GNC pipeline, a control-only experiment is run. The purpose of experimenting only with the controller is to efficiently evaluate a large variety of strategies and controller tuning. We study the use of DC versus the SLMP, mean state targeting via UT, and hysteresis via trigger condition and targeting tolerance tuning. The navigation error is drawn from a Gaussian distribution with position and velocity standard deviations of $3$-$\sigma_{r}=5$  $\text{\,}\mathrm{k}\mathrm{m}$ and $3\sigma_{v}=3$  $\text{\,}\mathrm{c}\mathrm{m}\mathrm{/}\mathrm{s}$; these values are chosen as conservative estimations that can be achieved with OPNAV based on the filter experiment with $f=360$  $\text{\,}\mathrm{m}\mathrm{m}$ and 3 measurements per revolution. For each controller configuration, 100 Monte-Carlo samples are taken, each lasting 60 NRHO revolutions, corresponding to about 393 days, using the same baseline NRHO as in Section 5. In all cases considered within this experiment, the SK maneuver is executed at a true anomaly of $180^{\circ}$. The various sources of error considered along with their 3-$\sigma$ values are given in Table 5.
@@ 6.1 Comparison of Differential Correction against SLMP
We first focus on the difference in performance between DC and SLMP, without and with hysteresis. Table 6 summarizes the statistics from the Monte-Carlo runs, and Figure 11 shows the distribution of the achieved violation of $v_{x}$ after control on the top row, and the yearly cost on the bottom row.
First, we observe by comparing cases 1A and 1B that the DC without or with hysteresis perform similarly; in contrast, looking at cases 1C and 1D, SLMP performs significantly worse than the other three cases when hysteresis is not used. The need for hysteresis is apparent by the top row in Figure 11, where the achieved $v_{x}$ violations between 1A and 1B are similar, and starkly different between 1C and 1D; in cases 1A, 1B and 1D a tight $v_{x}$ violation is achieved, while in 1C, the minimization problem aims to satisfy the targeting constraint at a higher tolerance, thus resulting in a higher cumulative cost.
Figure 11: Post trigger/control target violation and cumulative yearly cost with control schemes targeting $v_{x}$
@@ 6.2 Efficacy of UT-Mean State Targeting
We now compare the use of unscented transform. Table 7 summarizes the statistics from the Monte-Carlo runs. Figure 12 shows the distribution of the achieved violation of $v_{x}$ after control on the top row, and the yearly cost on the bottom row. Comparing 2A against 1B and 2B against 1D, we see that the use of UT improves the performance for both the DC and SLMP approach. This improvement validates the motivation of the UT-based approach: the effect of the errors on the initial state estimate is attenuated by incorporating sigma points to predict the final targeted state. The use of UT results in an increase in the average number of iterations required, as the problem is capturing the nonlinearity better, and both DC and SLMP are based on linearized correction schemes or dynamics, respectively.
Figure 12: Post trigger/control target violation and cumulative yearly cost with control schemes targeting $\mathbb{E}[v_{x}]$
7 GNC Pipeline Validation
The final set of numerical experiments consists of validating the entire pipeline, with online generation of synthetic images, simulation of onboard navigation filtering, and station-keeping. Building upon the insights from Section 5, we select a camera with $f=360$  $\text{\,}\mathrm{m}\mathrm{m}$, taking measurements at true anomalies of $145^{\circ}$, $155^{\circ}$, and $215^{\circ}$. We test controller configurations 1B and 2A, which were the cases with the best performance from the experiments in Section 6, without and with the UT-based targeting, respectively. In both cases, we select $v_{x,\mathrm{trig}}=20$  $\text{\,}\mathrm{m}\mathrm{/}\mathrm{s}$ and $v_{x,\mathrm{tol}}=1$  $\text{\,}\mathrm{m}\mathrm{/}\mathrm{s}$.
@@ 7.1 Comparison against Filter-Only and Controller-Only Experiments
Table 8 summarizes the yearly $\Delta V$, Figure 13 shows the distribution of the cumulative cost, and the corresponding cumulative cost histories are shown in Figure 14. The resulting $\Delta V$ are lower than those reported from the controller experiments in Tables 6 and 7, since the previously assumed $3$-$\sigma_{r}=5$  $\text{\,}\mathrm{k}\mathrm{m}$ and $3\sigma_{v}=3$  $\text{\,}\mathrm{c}\mathrm{m}\mathrm{/}\mathrm{s}$ were conservative. Figure 15 shows that the state estimate errors immediately before the maneuver are at lower values. Figure 16 shows the diagonal components of the covariance immediately before the maneuver, which are consistent with the empirical errors in Figure 15.
The relative performance between 1A and 2B is consistent both from Section 6 and the GNC pipeline results presented here, where we see a substantial benefit from UT-based targeting. From Table 8, the advantage of UT is particularly pronounced in the $95^{\mathrm{th}}$ percentile and max cost from the Monte-Carlo experiments; as shown in Figure 14, out of the 100 Monte-Carlo samples, controller 2A has a smaller standard deviation, and there are few samples where the error realizations result in higher cumulative cost with controller 1B. While the same realizations are also apparent for controller 2A, the cumulative cost does not accumulate as fast as in 1B.
Table 8: Summary of 100 Monte-Carlo runs comparing DC against UT-DC, using EKF with 3 measurements per revolution, $f=360$  $\text{\,}\mathrm{m}\mathrm{m}$ camera, $v_{x,\mathrm{trig}}=20$  $\text{\,}\mathrm{m}\mathrm{/}\mathrm{s}$ and $v_{x,\mathrm{tol}}=1$  $\text{\,}\mathrm{m}\mathrm{/}\mathrm{s}$, maneuvering at a true anomaly of $180^{\circ}$.
@@
Figure 13: Distribution of yearly station-keeping $\Delta V$ cost with full autonomous GNC pipeline, from 100 Monte-Carlo samples each; green triangle is the mean, and orange line is the median
Figure 14: Cumulative station-keeping cost with full autonomous GNC pipeline, from 100 Monte-Carlo samples each
Figure 15: Distribution of filter’s state estimate error immediately before maneuver at a true anomaly of $180^{\circ}$. Red dashed lines correspond to the empirical 22-$\sigma$ bounds.
Figure 16: Distribution of navigation filter’s standard deviation immediately before maneuver at a true anomaly of $180^{\circ}$
@@ 7.2 Sensitivity on Control Location
Guzzetti et al. [17] show that a constant covariance, for instance based on DSN measurements, results in a flat trend of total cumulative station-keeping cost against maneuver location. Using the full GNC pipeline consisting of the synthetic images and simulation of the navigation filters, we provide insight into the effect of the maneuver location with higher realism than assuming a fixed level of navigation error, as is common in control-only MC experiments that assume fixed navigation errors [16, 17, 19]. For this experiment, we adopt the UT-DC controller 2B, which was found to provide the best cost performance from results in 7.1.
Figure 17 shows the results from 100 Monte-Carlo runs, where the control action true anomaly is varied within a $\pm 15^{\circ}$ window about the apolune. While the average and median cost differences are below $10\text{\,}\mathrm{c}\mathrm{m}\mathrm{/}\mathrm{s}$, a larger discrepancy in the worst-case performance, exceeding a difference of $10\text{\,}\mathrm{c}\mathrm{m}\mathrm{/}\mathrm{s}$, is observed. Specifically, with respect to the considered NRHO, camera properties, and measurement frequency, maneuvering at a true anomaly of $170^{\circ}$ yields the best performance, both in terms of the mean/median, and worst-case cumulative costs. We can interpret the small, albeit non-negligible difference in these yearly cost distributions by looking at the filter standard deviations from Figure 8. The minimum in $\sigma_{v_{z}}$ occurs at around $170^{\circ}$, which coincides with the best-performing maneuver location in Figure 17. Such an observation is made possible by the realistic assessment that couples the navigation filter performance with station-keeping, especially in cases where a constant navigation covariance is not applicable.
Figure 17: Distribution of navigation filter’s standard deviation immediately before maneuver; green triangle is the mean, and orange line is the median
Table 9: Summary of 100 Monte-Carlo runs on station-keeping cost with UT-DC controller, executing maneuver at varying true anomalies around apolune. All cases using EKF with 3 measurements per revolution, $f=360$  $\text{\,}\mathrm{m}\mathrm{m}$ camera, $v_{x,\mathrm{trig}}=20$  $\text{\,}\mathrm{m}\mathrm{/}\mathrm{s}$ and $v_{x,\mathrm{tol}}=1$  $\text{\,}\mathrm{m}\mathrm{/}\mathrm{s}$.
@@
165 100% 43.7387 56.2019 69.0637
170 100% 43.7115 58.0249 63.6931
175 100% 44.6368 59.8371 69.7114
180 100% 47.9092 62.5284 76.1859
185 100% 49.7340 68.6846 80.0424
190 100% 50.0258 68.6872 73.8191
195 100% 51.0217 64.5657 79.0357
@@ 8 Conclusion
Several key insights have been obtained from the development and evaluation of the guidance, navigation, and control pipeline for operations in a Near Rectilinear Halo Orbit with horizon-based OPNAV.
From a navigation standpoint, the simulations demonstrate that only a limited number of horizon-based measurements are necessary for the filter to maintain sufficient state estimation accuracy on the NRHO. An investigation of the camera field of view and measurement locations along the orbit reveals notable variations in estimation error. Across all tested configurations, the filter covariance exhibits a persistent and well-defined periodic structure. The structure is attributed to the inherent degradation of horizon-based measurement precision with increasing true anomaly, driven primarily by the rapid reduction in the apparent diameter of the Moon.
On the control side, the quasi-Newton convergence properties of differential correction provide an advantage in reducing cumulative station-keeping cost compared to an explicit minimization-based strategy. The lower performance of the latter arises from its short-sighted per-maneuver cost objective that only satisfies the targeting constraint to a specified tolerance, which does not directly minimize the overall cumulative cost. The introduction of hysteresis improves the performance for both differential correction and minimization-based approaches, as choosing a smaller targeting tolerance than the triggering tolerance enhances robustness against maneuver execution error and model uncertainty. Additionally, incorporating an unscented transform-based targeting framework substantially reduces the $95^{\mathrm{th}}$-percentile and worst-case station-keeping costs observed in Monte Carlo simulations. By leveraging the filter covariance rather than relying solely on the state estimate, the unscented transform-based framework mitigates sensitivity to estimation errors, which vary considerably across realizations. Finally, the adoption of OPNAV-based navigation introduces a periodic structure in the filter covariance, which directly influences station-keeping performance, as the cumulative station-keeping cost becomes dependent on the maneuver location along the NRHO.
The OPNAV-based GNC pipeline developed in this work can be extended to a broader class of libration point orbits for which horizon-based OPNAV is applicable. Beyond horizon-based techniques, alternative OPNAV methods such as crater-based navigation and visual odometry may be considered to further enhance autonomous onboard navigation and reduce reliance on ground-based updates.

### v2

@@ 6 Controller Experiments
Prior to running the full GN&C pipeline, a control-only experiment is run. The purpose of experimenting only with the controller is to efficiently evaluate a large variety of strategies and controller tuning. We study the use of DC versus the SLMP, mean state targeting via UT, and hysteresis via trigger condition and targeting tolerance tuning. The navigation error is drawn from a Gaussian distribution with position and velocity standard deviations of $3$-$\sigma_{r}=5$  $\text{\,}\mathrm{k}\mathrm{m}$ and $3\sigma_{v}=3$  $\text{\,}\mathrm{c}\mathrm{m}\mathrm{/}\mathrm{s}$; these values are chosen as conservative estimations that can be achieved with OPNAV based on the filter experiment with $f=360$  $\text{\,}\mathrm{m}\mathrm{m}$ and 3 measurements per revolution. To focus on the controller’s tracking behavior, Monte-Carlo experiments in this Section do not include any initial insertion error. For each controller configuration, 100 Monte-Carlo samples are taken, each lasting 60 NRHO revolutions, corresponding to about 393 days, using the same baseline NRHO as in Section 5. In all cases considered within this experiment, the SK maneuver is executed at $\theta=180^{\circ}$. The various sources of error considered along with their 3-$\sigma$ values are given in Table 4. In this work, we assume a “quiet” spacecraft station-keeping scenario [Davis2017], where no random impulses are imparted due to attitude control.
@@ 6.1 Comparison of Differential Correction against SLMP
We first focus on the difference in performance between DC and SLMP, without and with hysteresis. Table 5 summarizes the statistics from the Monte-Carlo runs, and Figure 10(a) shows the distribution of the achieved violation of $v_{x}$ after control on the top row, and the yearly cost on the bottom row.
We observe by comparing cases 1A and 1B that the DC without or with hysteresis performs similarly; in contrast, looking at cases 1C and 1D, SLMP performs significantly worse than the other three cases when hysteresis is not used. The need for hysteresis is apparent by the top row in Figure 10(a), where the achieved $v_{x}$ violations between 1A and 1B are similar, and starkly different between 1C and 1D; in cases 1A, 1B and 1D a tight $v_{x}$ violation is achieved, while in 1C, the minimization problem aims to satisfy the targeting constraint at a higher tolerance, thus resulting in a higher cumulative cost. With SLMP, a tighter $v_{x,\rm tol}$ reduces the yearly station-keeping cost, highlighting that the long-term benefit of aligning the controlled trajectory closer to the reference outweighs the higher maneuver cost for a maneuver instance. Such an effect is not pronounced for DC due to its quadratic convergence of $|v_{x}^{\rm EM}-\bar{v}_{x}^{\rm EM}|\ll v_{x,\rm tol}$. In cases within 1A and 1B, we find similar performances with $v_{x,\rm trig}=10$  $\text{\,}\mathrm{m}\mathrm{/}\mathrm{s}$ and $v_{x,\rm trig}=20$  $\text{\,}\mathrm{m}\mathrm{/}\mathrm{s}$, and a comparatively worse performance with $v_{x,\rm trig}=30$  $\text{\,}\mathrm{m}\mathrm{/}\mathrm{s}$. Such an increase in cumulative cost is attributed to the impact of an excessively high $v_{x,\rm trig}$, which results in less frequent, but cumulatively costly station-keeping controls.
(a) $v_{x}$-targeting schemes
(b) $\mathbb{E}[v_{x}]$-targeting schemes
Figure 10: Post-trigger/post-control $v_{x}$ violation and yearly cost
We verify that the controlled trajectory follows the reference NRHO by looking at its position history. Figure 11 shows the controlled trajectory over 60 revolutions in $\mathcal{F}_{\rm EM}$ for a single sample out of the Monte-Carlo runs from 1B with $v_{x,\rm trig}=20$  $\text{\,}\mathrm{m}\mathrm{/}\mathrm{s}$. Controlled trajectories from other samples, as well as other controllers, closely resemble Figure 11. We also show the norm of the position vector deviation over time, $\Delta r(t)=\|\boldsymbol{r}(t)-\bar{\boldsymbol{r}}(t)\|_{2}$, from 1B in Figure 12. We observe a slow secular growth, but that remains under $100$  $\text{\,}\mathrm{k}\mathrm{m}$ over the course of 60 revolutions. Comparing with Figure 11, the growth in $\Delta r$ is primarily due to along-track deviation, as the qualitative geometry of the NRHO is maintained throughout the 60 revolutions. The $\Delta r$ histories from other controllers also closely resemble Figure 12.
The secular growth in along-track deviation is a known side-effect of $x$-axis crossing control, and is related to its reliance on event-based targeting [Shimane2024-ep], which prioritizes maintaining the geometry of the orbit but neglects epoch deviation. As a remedy, an additional infrequent (e.g., once a year) targeting maneuver may be executed to realign the trajectory to the reference [Davis2022], or the $x$-axis crossing control may be replaced by a controller that recursively tracks the full-state [Shimane2025].
Figure 11: Example controlled trajectory over 60 revolutions from 1B
Figure 12: Position deviation histories from 1B
@@ 6.2 Efficacy of UT-Mean State Targeting
We now compare the use of the unscented transform. Table 6 summarizes the statistics from the Monte-Carlo runs. Figure 10(b) shows the distribution of the achieved violation of $v_{x}$ after control on the top row, and the yearly cost on the bottom row. Comparing 2A against 1B and 2B against 1D, we see that the use of UT improves the performance for both the DC and SLMP approach. This improvement validates the motivation of the UT-based approach: the effect of the errors on the initial state estimate is attenuated by incorporating sigma points to predict the final targeted state. The use of UT results in an increase in the average number of iterations required, as the problem is capturing the nonlinearity better, and both DC and SLMP are based on linearized correction schemes or dynamics, respectively. The controlled trajectories and trend in $\Delta r$ histories resembles those shown in Figures 11 and 12, and are omitted for conciseness.
7 GN&C Pipeline Validation
The final set of numerical experiments consists of validating the entire pipeline, with online generation of synthetic images, simulation of onboard navigation filtering, and station-keeping. Building upon the insights from Section 5, we select a camera with $f=360$  $\text{\,}\mathrm{m}\mathrm{m}$, taking measurements at true anomalies of $145^{\circ}$, $155^{\circ}$, and $215^{\circ}$. We test controller configurations 1B and 2A, which were the cases with the best performance from the experiments in Section 6, without and with the UT-based targeting, respectively. In both cases, we select $v_{x,\mathrm{trig}}=20$  $\text{\,}\mathrm{m}\mathrm{/}\mathrm{s}$ and $v_{x,\mathrm{tol}}=1$  $\text{\,}\mathrm{m}\mathrm{/}\mathrm{s}$. In each experiment in this Section, we consider 100 Monte-Carlo runs. As with Monte-Carlo runs from Sections, we re-introduce a random initial insertion error, with standard deviations $3\text{-}\sigma_{r}=10$  $\text{\,}\mathrm{k}\mathrm{m}$ and $3\text{-}\sigma_{v}=10$  $\text{\,}\mathrm{c}\mathrm{m}\mathrm{/}\mathrm{s}$.
@@ 7.1 Comparison against Filter-Only and Controller-Only Experiments
Table 7 summarizes the yearly $\Delta V$, Figure 13 shows the distribution of the cumulative cost, and the corresponding cumulative cost histories are shown in Figure 14. The resulting $\Delta V$ are lower than those reported from the controller experiments in Tables 5 and 6, since the previously assumed $3$-$\sigma_{r}=5$  $\text{\,}\mathrm{k}\mathrm{m}$ and $3\sigma_{v}=3$  $\text{\,}\mathrm{c}\mathrm{m}\mathrm{/}\mathrm{s}$ were conservative. Figure 15 shows that the state estimate errors immediately before the maneuver are at lower values. Figure 16 shows the diagonal components of the covariance immediately before the maneuver, which are consistent with the empirical errors in Figure 15.
The relative performance between 1A and 2B is consistent both from Section 6 and the GN&C pipeline results presented here, where we see a substantial benefit from UT-based targeting. From Table 7, the advantage of UT is particularly pronounced in the $95^{\mathrm{th}}$ percentile and max cost from the Monte-Carlo experiments; as shown in Figure 14, out of the 100 Monte-Carlo samples, controller 2A has a smaller standard deviation, and there are few samples where the error realizations result in higher cumulative cost with controller 1B. While the same realizations are also apparent for controller 2A, the cumulative cost does not accumulate as fast as in 1B.
Table 7: Monte-Carlo summary from DC and UT-DC, using EKF with 3 measurements per revolution, $f=360$  $\text{\,}\mathrm{m}\mathrm{m}$ camera, $v_{x,\mathrm{trig}}=20$  $\text{\,}\mathrm{m}\mathrm{/}\mathrm{s}$ and $v_{x,\mathrm{tol}}=1$  $\text{\,}\mathrm{m}\mathrm{/}\mathrm{s}$, maneuvering at $\theta=180^{\circ}$
@@
Figure 13: Distribution of yearly cost with full autonomous GN&C pipeline
(a) 1B ($v_{x,\rm trig}=20$  $\text{\,}\mathrm{m}\mathrm{/}\mathrm{s}$, $v_{x,\rm tol}=1$  $\text{\,}\mathrm{m}\mathrm{/}\mathrm{s}$)
(b) 2A ($v_{x,\rm trig}=20$  $\text{\,}\mathrm{m}\mathrm{/}\mathrm{s}$, $v_{x,\rm tol}=1$  $\text{\,}\mathrm{m}\mathrm{/}\mathrm{s}$)
Figure 14: Cumulative station-keeping cost with full autonomous GN&C pipeline
Figure 15: Distribution of filter’s state estimate error immediately before maneuver at $\theta=180^{\circ}$
Figure 16: Navigation filter’s standard deviation distribution immediately before maneuver at $\theta=180^{\circ}$
@@ 7.2 Sensitivity on Control Location
Guzzetti et al. [Guzzetti2017] show that a constant covariance, for instance based on DSN measurements, results in a flat trend of total cumulative station-keeping cost against maneuver location. Using the full GN&C pipeline consisting of the synthetic images and simulation of the navigation filters, we provide insight into the effect of the maneuver location with higher realism than assuming a fixed level of navigation error, as is common in control-only MC experiments that assume fixed navigation errors [Davis2017, Guzzetti2017, Davis2022]. For this experiment, we adopt the UT-DC controller 2B, which was found to provide the best cost performance from results in 7.1.
Figure 17 shows the results from 100 Monte-Carlo runs with 2B, where the control action true anomaly is varied within a $\pm 15^{\circ}$ window about the apolune. While the average and median cost differences are below $10\text{\,}\mathrm{c}\mathrm{m}\mathrm{/}\mathrm{s}$, a larger discrepancy in the worst-case performance, exceeding a difference of $10\text{\,}\mathrm{c}\mathrm{m}\mathrm{/}\mathrm{s}$, is observed. Specifically, with respect to the considered NRHO, camera properties, and measurement frequency, maneuvering at a true anomaly of $170^{\circ}$ yields the best performance, both in terms of the mean/median, and worst-case cumulative costs. We can interpret the small, albeit non-negligible difference in these yearly cost distributions by looking at the filter standard deviations from Figure 9(b). The minimum in $\sigma_{v_{z}}$ occurs at around $170^{\circ}$, which coincides with the best-performing maneuver location in Figure 17. Such an observation is made possible by the realistic assessment that couples the navigation filter performance with station-keeping, especially in cases where a constant navigation covariance is not applicable.
Figure 17: Distribution of yearly cost with UT-DC controller across varying maneuver true anomaly
Table 8: Station-keeping cost with UT-DC controller for varying maneuver true anomaly
@@
$165^{\circ}$ 100% 43.7387 56.2019 69.0637
$170^{\circ}$ 100% 43.7115 58.0249 63.6931
$175^{\circ}$ 100% 44.6368 59.8371 69.7114
$180^{\circ}$ 100% 47.9092 62.5284 76.1859
$185^{\circ}$ 100% 49.7340 68.6846 80.0424
$190^{\circ}$ 100% 50.0258 68.6872 73.8191
$195^{\circ}$ 100% 51.0217 64.5657 79.0357
@@ 8 Conclusion
Several key insights have been obtained from the development and evaluation of the guidance, navigation, and control pipeline for operations in a Near Rectilinear Halo Orbit with horizon-based OPNAV. From a navigation standpoint, the simulations demonstrate that only a limited number of horizon-based measurements are necessary for the filter to maintain sufficient state estimation accuracy on the NRHO. An investigation of the camera field of view and measurement locations along the orbit reveals notable variations in estimation error. Across all tested configurations, the filter covariance exhibits a persistent and well-defined periodic structure. The structure is attributed to the inherent degradation of horizon-based measurement precision with increasing true anomaly, driven primarily by the rapid reduction in the apparent diameter of the Moon. With the EKF using 3 measurements collected per revolution, we find a consistent global minimum in standard deviations at a true anomaly of around $170^{\circ}$.
On the control side, the quasi-Newton convergence properties of differential correction provide an advantage in reducing cumulative station-keeping cost compared to an explicit minimization-based strategy. The lower performance of the latter, arises from its short-sighted per-maneuver cost objective that only satisfies the targeting constraint to a specified tolerance, which does not directly minimize the overall cumulative cost. The introduction of hysteresis improves the performance for both differential correction and minimization-based approaches, as choosing a smaller targeting tolerance than the triggering tolerance enhances robustness against maneuver execution error and model uncertainty. Additionally, incorporating an unscented transform-based targeting framework substantially reduces the $95^{\mathrm{th}}$-percentile and worst-case station-keeping costs observed in Monte Carlo simulations. By leveraging the filter covariance rather than relying solely on the state estimate, the unscented transform-based framework mitigates sensitivity to estimation errors, which vary considerably across realizations. Finally, the adoption of OPNAV-based navigation introduces a periodic structure in the filter covariance, which directly influences station-keeping performance, as the cumulative station-keeping cost becomes dependent on the maneuver location along the NRHO. Using the UT-DC controller with maneuvers executed at the global minimum of the EKF standard deviation, we successfully maintain the NRHO at a yearly station-keeping cost of about $60\text{\,}\mathrm{c}\mathrm{m}\mathrm{/}\mathrm{s}$ at the $95^{\rm th}$ percentile.
The OPNAV-based GN&C pipeline developed in this work can be extended to a broader class of libration point orbits for which horizon-based OPNAV is applicable. Beyond horizon-based techniques, alternative OPNAV methods such as crater-based navigation and visual odometry may be considered to further enhance autonomous onboard navigation and reduce reliance on ground-based updates.